OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to y = 3x+6 and passes through the point (-10, 2.5)

OpenStudy (anonymous):

(y-2.5)=3(x+10) y-2.5=3x+30 y=3x+32.5